# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ C.3  Auxilary lemmas for the oracle inequality
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ is a sequence of i.i.d. Rademacher variables. Here the first inequality follows from Lemma C.10, and the second inequality is due to the fact that $\widehat{h}_{t}^{\lambda}\in{\mathscr{H}}_{t}$, the second inequality holds from Lemma C.10. This completes the proof. ∎
C.4  Convergence rate of the Doob’s guidance estimation
Lemma C.12.
@@ C.4  Convergence rate of the Doob’s guidance estimation
Suppose Assumptions 1 and 2 hold. Let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (4.3) with $L\leq C\log N$ and $S\leq N^{d}$, then there exists $h_{t}\in{\mathscr{H}}_{t}$ such that
$\|h_{t}-h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $\leq\frac{CR^{2}}{\sigma_{T-t}^{4}N^{2}},$
$\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $\leq\frac{CR}{\sigma_{T-t}^{4}N},$
where $C$ is a constant independent of $N$ and $R$.
Proof of Lemma C.12.
@@
We first rescale the target function $h_{t}^{*}$ to $B({\mathbf{0}},1)$ by $g_{t}^{*}(\mathbf{z})\coloneq h_{t}^{*}(R\mathbf{z})$. According to Ding et al. [2025, Lemma 6], there exists $g_{t}\in N(L,S)$ such that
$\|g_{t}-g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}$ $\leq\frac{C}{N^{2}}\|g_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
$\|\nabla g_{t}-\nabla g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}$ $\leq\frac{C}{N}\|g_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
@@
where $C$ is a constant only depend on $d$. Note that $D_{k}g_{t}^{*}(\mathbf{z})=RD_{k}h_{t}^{*}(R\mathbf{z})$ for each $1\leq k\leq d$, and $D_{k\ell}^{2}g_{t}^{*}(\mathbf{z})=R^{2}D_{k\ell}^{2}h_{t}^{*}(R\mathbf{z})$ for each $1\leq k,\ell\leq d$. Thus
$\|g_{t}(R^{-1}\cdot)-h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $=\|g_{t}(R^{-1}\cdot)-g_{t}^{*}(R^{-1}\cdot)\|_{L^{\infty}(B({\mathbf{0}},1))}\leq\frac{CR^{2}}{N^{2}}\|h_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
$\|\nabla g_{t}(R^{-1}\cdot)-\nabla h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $=\frac{1}{R}\|\nabla g_{t}-\nabla g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}\leq\frac{CR}{N}\|h_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})}.$
Set $h_{t}\coloneq g_{t}(R^{-1}\cdot)$, and using Lemmas C.1, C.4, and C.5 complete the proof. ∎
Lemma C.13 (Approximation error).
@@
Suppose Assumptions 1 and 2 hold. Let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (4.3) with $L\leq C\log N$ and $S\leq N^{d}$, then
$\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{CR^{4}}{\sigma_{T-t}^{8}N^{2}}+\frac{C}{\sigma_{T-t}^{4}}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where $C$ is a constant independent of $t$, NN and $R$. Further, by setting $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{2}$, we have
$\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{C^{\prime}\log^{2}N}{\sigma_{T-t}^{8}N^{2}},$
where $C^{\prime}$ is a constant independent of $t$ and $N$.
Proof of Lemma C.13.
@@
(C.23) $\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}$
$=\underbrace{\int\Big((h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}+\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}\Big){\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(i)}}$
$\quad+\underbrace{\int\Big((h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}+\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}\Big){\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(ii)}}.$
For term (i) in (C.23), we have
$\int\Big((h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}+\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}\Big){\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(C.24) $\leq\sup_{\|\mathbf{x}\|_{2}\leq R}\Big((h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}+\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}\Big)\leq\frac{CR^{4}}{\sigma_{T-t}^{8}N^{2}},$
where the second inequality holds from Lemma C.12. For term (ii) in (C.23), we have
@@
$\leq 4\Big(\bar{B}^{2}+\frac{4\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big)\operatorname{Pr}\{\|\mathbf{X}_{T-t}\|_{2}>R\}$
(C.25) $\leq 2^{d+3}\Big(\bar{B}^{2}+\frac{4\bar{B}^{2}}{\sigma_{T-t}^{4}}\Big)\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where the first inequality holds from Lemmas C.1 and C.4, and the second inequality is due to Lemma E.1. Combining (C.23), (C.24), and (C.25) yields
$\inf_{h_{t}\in{\mathscr{H}}_{t}}\|h_{t}-h_{t}^{*}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{CR^{4}}{\sigma_{T-t}^{8}N^{2}}+\frac{C}{\sigma_{T-t}^{4}}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Setting $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{2}$ completes the proof. ∎
Lemma C.14 (Generalization error).
@@
${\mathfrak{R}}_{n}({\mathscr{H}}_{t})$ $\leq c\bar{B}\Big(\frac{LS\log(S)}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
${\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})$ $\leq c\frac{\bar{B}}{\sigma_{T-t}^{2}}\Big(\frac{L^{2}S\log(LS)}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
@@
Proof of Lemma C.14.
Define a function class $\widetilde{{\mathscr{H}}}_{t}$ as
(C.26) $\widetilde{{\mathscr{H}}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}.\end{aligned}\right\}.$
It is apparent that ${\mathscr{H}}_{t}=\mathrm{conv}(\widetilde{{\mathscr{H}}}_{t})$, and $\nabla{\mathscr{H}}_{t}=\mathrm{conv}(\nabla\widetilde{{\mathscr{H}}}_{t})$. Then applying Bartlett and Mendelson [2002, Theorem 12] implies
(C.27) ${\mathfrak{R}}({\mathscr{H}}_{t}|\mathbf{X}_{T-t}^{1:n})$ $={\mathfrak{R}}(\widetilde{{\mathscr{H}}}_{t}|\mathbf{X}_{T-t}^{1:n}),$
(C.28) ${\mathfrak{R}}(\nabla{\mathscr{H}}_{t}|\mathbf{X}_{T-t}^{1:n})$ $={\mathfrak{R}}(\nabla\widetilde{{\mathscr{H}}}_{t}|\mathbf{X}_{T-t}^{1:n}),$
where $\mathbf{X}_{T-t}^{1},\ldots,\mathbf{X}_{T-t}^{n}\sim^{\mathrm{i.i.d.}}p_{T-t}$. Let $\delta>0$ and $\widetilde{{\mathscr{H}}}_{t}^{\delta}$ be an $L^{\infty}(\mathbf{X}_{T-t}^{1:n})$ $\delta$-cover of $\widetilde{{\mathscr{H}}}_{t}$ satisfying $|\widetilde{{\mathscr{H}}}_{t}^{\delta}|=N(\delta,\widetilde{{\mathscr{H}}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))$. Then for each $h_{t}\in\widetilde{{\mathscr{H}}}_{t}$, there exists $h_{t}^{\delta}\in\widetilde{{\mathscr{H}}}_{t}^{\delta}$ such that
As a consequence,
${\mathfrak{R}}(\widetilde{{\mathscr{H}}}_{t}|\mathbf{X}_{T-t}^{1:n})$ $={\mathbb{E}}\Bigg[\sup_{h_{t}\in\widetilde{{\mathscr{H}}}_{t}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}(\mathbf{X}_{t}^{i})\Bigg|\mathbf{X}_{T-t}^{1:n}\Bigg]$
$\leq{\mathbb{E}}\Bigg[\sup_{h_{t}^{\delta}\in\widetilde{{\mathscr{H}}}_{t}^{\delta}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h_{t}^{\delta}(\mathbf{X}_{t}^{i})\Bigg|\mathbf{X}_{T-t}^{1:n}\Bigg]+\delta$
$\leq\bar{B}\Big(\frac{2\log|\widetilde{{\mathscr{H}}}_{t}^{\delta}|}{n}\Big)^{\frac{1}{2}}+\delta$
$=\bar{B}\Big(\frac{2\log N(\delta,\widetilde{{\mathscr{H}}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))}{n}\Big)^{\frac{1}{2}}+\delta,$
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ is a sequence of i.i.d. Rademacher variables, the second inequality follows from Massart’s lemma [Mohri et al., 2018, Theorem 3.7]. Setting $\delta=\bar{B}/\sqrt{n}$ yields
(C.29) ${\mathfrak{R}}(\widetilde{{\mathscr{H}}}_{t}|\mathbf{X}_{T-t}^{1:n})\leq\bar{B}\Big(\frac{2\log N(B/\sqrt{n},\widetilde{{\mathscr{H}}}_{t},L^{\infty}(\mathbf{X}_{T-t}^{1:n}))}{n}\Big)^{\frac{1}{2}}\leq 8\bar{B}\Big(\frac{\mathrm{VCdim}(\widetilde{{\mathscr{H}}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where the last inequality holds from Anthony et al. [1999, Theorem 12.2]. Since $\widetilde{{\mathscr{H}}}_{t}\subseteq N(L,S)$, using Bartlett et al. [2019, Theorem 7] implies
(C.30) $\mathrm{VCdim}(\widetilde{{\mathscr{H}}}_{t})\leq\mathrm{VCdim}(N(L,S))\leq c_{0}LS\log(S),$
where $c_{0}$ is an absolute constant. Combining (C.27), (C.29), and (C.30) deduces
${\mathfrak{R}}_{n}({\mathscr{H}}_{t})={\mathbb{E}}[{\mathfrak{R}}({\mathscr{H}}_{t}|\mathbf{X}_{T-t}^{1:n})]\leq c_{1}\bar{B}\Big(\frac{LS\log(S)}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where $c_{1}$ is an absolute constant. By a similar argument as (C.29), we have
(C.31) ${\mathfrak{R}}(\nabla\widetilde{{\mathscr{H}}}_{t}|\mathbf{X}_{T-t}^{1:n})\leq 16\frac{\bar{B}}{\sigma_{T-t}^{2}}\Big(\frac{\mathrm{VCdim}(\nabla\widetilde{{\mathscr{H}}}_{t})}{n\log^{-1}n}\Big)^{\frac{1}{2}}.$
Using, we have $\nabla\widetilde{{\mathscr{H}}}_{t}\subseteq N(c_{2}L,c_{3}LS)$, where $c_{2}$ and $c_{3}$ are absolute constants. Using Bartlett et al. [2019, Theorem 7] implies
(C.32) $\mathrm{VCdim}(\nabla\widetilde{{\mathscr{H}}}_{t})\leq c_{4}L^{2}S\log(LS),$
where $c_{4}$ is an absolute constant. Combining (C.28), (C.31), and (C.32) deduces
${\mathfrak{R}}_{n}(\nabla{\mathscr{H}}_{t})={\mathbb{E}}[{\mathfrak{R}}(\nabla{\mathscr{H}}_{t}|\mathbf{X}_{T-t}^{1:n})]\leq c_{5}\frac{\bar{B}}{\sigma_{T-t}^{2}}\Big(\frac{L^{2}S\log(LS)}{n\log^{-1}n}\Big)^{\frac{1}{2}},$
where $c_{5}$ is an absolute constant. This completes the proof. ∎
@@
(C.33) ${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in\mathrm{conv}(N(L,S)):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B},\end{aligned}\right\}.$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+4}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient penalized empirical risk minimizer defined as (3.21), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequalities hold:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{1}{d+4}},$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{1}{d+4}})$, where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
@@
Substituting Lemmas C.13, and C.14 into Lemma 4.2 yields
@@
$\lesssim\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+\Big(\frac{LS\log(S)}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda}{\sigma_{T-t}^{4}}\Big(\frac{L^{2}S\log(LS)}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda^{2}}{\sigma_{T-t}^{8}}$
$\lesssim\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+\Big(\frac{N^{d}\log^{2}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda}{\sigma_{T-t}^{4}}\Big(\frac{N^{d}\log^{4}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+\frac{\lambda^{2}}{\sigma_{T-t}^{8}},$
where we omit the constants independent of $t$, NN and $n$, and the second inequality is due to $L\leq C\log N$ and $S\leq N^{d}$. Similarly,

### v2

@@ D.2  Auxilary lemmas for the oracle inequality
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ is a sequence of i.i.d. Rademacher variables. Here the first inequality follows from Lemma D.4, and the second inequality is due to the fact that $\widehat{h}_{t}^{\lambda}\in{\mathscr{H}}_{t}$, the second inequality holds from Lemma D.4. Finally, using (D.11) completes the proof. ∎
Lemma D.6.
Suppose Assumptions 1 and 2 hold. Then
(D.12) $\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}\leq\frac{12\sqrt{d}\bar{B}}{\sigma_{T-t}^{4}}.$
Proof of Lemma D.6.
By applying the triangular inequality, we have
$\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}$
$\leq\|\Delta h_{t}^{*}\|_{L^{2}(p_{T-t})}+\|\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}$
(D.13) $\leq\frac{6\sqrt{d}\bar{B}}{\sigma_{T-t}^{4}}+\frac{2\bar{B}}{\sigma_{T-t}^{2}}\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})},$
where the last inequality holds from Lemmas C.4 and C.5. It remains to estimate the $L^{2}(p_{T-t})$-norm of the score $\nabla\log p_{T-t}$ in (D.13). Indeed,
$\|\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}$
(D.14) $=\frac{2}{\sigma_{T-t}^{4}}\Big\{\mu_{T-t}^{2}{\mathbb{E}}\big[\|\mathbf{X}_{0}\|_{2}^{2}\big]+\sigma_{T-t}^{2}{\mathbb{E}}\big[\|{\bm{\varepsilon}}\|_{2}^{2}\big]\Big\}+\frac{2d\mu_{T-t}^{2}}{\sigma_{T-t}^{4}}\leq\frac{6d}{\sigma_{T-t}^{4}},$
where the first equality is owing to Lemma C.2, the first inequality used Assumption 1, and the second and third equalities hold from $\mathbf{X}_{T-t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}{\bm{\varepsilon}}$ where $\mathbf{X}_{0}$ is independent ${\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})$. The last inequality also uses Assumption 1. Substituting (D.14) into (D.13) completes the proof. ∎
D.3  Convergence rate of the Doob’s guidance estimation
Lemma D.7.
@@ D.3  Convergence rate of the Doob’s guidance estimation
Suppose Assumptions 1 and 2 hold. Let $R\geq 1$, and let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.2) with $L\leq C\log N$ and $S\leq N^{d}$, then there exists $h_{t}\in{\mathscr{H}}_{t}$ such that
$\|h_{t}-h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $\leq\frac{C\bar{B}R^{2}}{\sigma_{T-t}^{4}N^{2}},$
$\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $\leq\frac{C\bar{B}R}{\sigma_{T-t}^{4}N},$
where $C$ is a constant only depending on $d$.
Proof of Lemma D.7.
@@
We first rescale the target function $h_{t}^{*}$ to $B({\mathbf{0}},1)$ by $g_{t}^{*}(\mathbf{z})\coloneq h_{t}^{*}(R\mathbf{z})$. According to Ding et al. [2025b, Lemma 6], there exists $g_{t}\in N(L,S)$ such that
$\|g_{t}-g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}$ $\leq\frac{C^{\prime}}{N^{2}}\|g_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
$\|\nabla g_{t}-\nabla g_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},1))}$ $\leq\frac{C^{\prime}}{N}\|g_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
@@
where $C^{\prime}$ is a constant only depending on $d$. Note that $D_{k}g_{t}^{*}(\mathbf{z})=RD_{k}h_{t}^{*}(R\mathbf{z})$ for each $1\leq k\leq d$, and $D_{k\ell}^{2}g_{t}^{*}(\mathbf{z})=R^{2}D_{k\ell}^{2}h_{t}^{*}(R\mathbf{z})$ for each $1\leq k,\ell\leq d$. Thus
$\|g_{t}(R^{-1}\cdot)-h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $=\|g_{t}(R^{-1}\cdot)-g_{t}^{*}(R^{-1}\cdot)\|_{L^{\infty}(B({\mathbf{0}},1))}\leq\frac{C^{\prime}R^{2}}{N^{2}}\|h_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})},$
$\|\nabla g_{t}(R^{-1}\cdot)-\nabla h_{t}^{*}\|_{L^{\infty}(B({\mathbf{0}},R))}$ $=\frac{1}{R}\|\nabla g_{t}(R^{-1}\cdot)-\nabla g_{t}^{*}(R^{-1}\cdot)\|_{L^{\infty}(B({\mathbf{0}},1))}\leq\frac{C^{\prime}R}{N}\|h_{t}^{*}\|_{C^{2}({\mathbb{R}}^{d})}.$
Setting $h_{t}\coloneq g_{t}(R^{-1}\cdot)$, and using Lemmas C.1, C.4, and C.5 complete the proof. ∎
Lemma D.8 (Approximation error).
@@
Suppose Assumptions 1 and 2 hold. Let $R\geq 1$, and let the hypothesis class ${\mathscr{H}}_{t}$ be defined as (5.2) with $L\leq C\log N$ and $S\leq N^{d}$, then
$\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{4}N}{\sigma_{T-t}^{8}N^{4}},$
$\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq C\frac{\bar{B}^{2}\log^{2}N}{\sigma_{T-t}^{8}N^{2}}.$
provided that $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{4}$, where $C$ is a constant only depending on $d$.
Proof of Lemma D.8.
@@
(D.15) $\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $=\underbrace{\int(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(i)}}$
$\quad+\underbrace{\int(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(ii)}}.$
For term (i) in (D.15), we have
$\int(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.16) $\leq\sup_{\|\mathbf{x}\|_{2}\leq R}(h_{t}(\mathbf{x})-h_{t}^{*}(\mathbf{x}))^{2}\leq\frac{C^{2}\bar{B}^{2}R^{4}}{\sigma_{T-t}^{8}N^{4}},$
where the second inequality holds from Lemma D.7. For term (ii) in (D.15), we have
@@
(D.17) $\leq 4\bar{B}^{2}\operatorname{Pr}\{\|\mathbf{X}_{T-t}\|_{2}>R\}\leq 2^{d+3}\bar{B}^{2}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where the first inequality holds from Lemma C.1, and the second inequality is due to Lemma G.1. Substituting (D.16) and (D.17) into (D.15) yields
(D.18) $\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\frac{C^{2}\bar{B}^{2}R^{4}}{\sigma_{T-t}^{8}N^{4}}+2^{d+3}\bar{B}^{2}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Similarly, for the gradient term, we have
(D.19) $\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $=\underbrace{\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(i)}}$
$\quad+\underbrace{\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}}_{\text{(ii)}}.$
For term (i) in (D.19), we have
$\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}\leq R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.20) $\leq\sup_{\|\mathbf{x}\|_{2}\leq R}\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}\leq\frac{C^{2}\bar{B}^{2}R^{2}}{\sigma_{T-t}^{8}N^{2}},$
where the second inequality holds from Lemma D.7. For term (ii) in (D.19), we have
$\int\|\nabla h_{t}(\mathbf{x})-\nabla h_{t}^{*}(\mathbf{x})\|_{2}^{2}{\mathbbm{1}}\{\|\mathbf{x}\|_{2}>R\}p_{T-t}(\mathbf{x})\,\mathrm{d}\mathbf{x}$
(D.21) $\leq\frac{16\bar{B}^{2}}{\sigma_{T-t}^{4}}\operatorname{Pr}\{\|\mathbf{X}_{T-t}\|_{2}>R\}\leq 2^{d+5}\frac{\bar{B}^{2}}{\sigma_{T-t}^{4}}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big),$
where the first inequality holds from Lemma C.4, and the second inequality is due to Lemma G.1. Substituting (D.20) and (D.21) into (D.19) yields
(D.22) $\|\nabla h_{t}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\leq\frac{C^{2}\bar{B}^{2}R^{2}}{\sigma_{T-t}^{8}N^{2}}+2^{d+5}\frac{\bar{B}^{2}}{\sigma_{T-t}^{4}}\exp\Big(-\frac{R^{2}}{4d\mu_{t}^{2}+8\sigma_{t}^{2}}\Big).$
Setting $R^{2}=(4d\mu_{t}^{2}+8\sigma_{t}^{2})\log N^{4}$ in (D.18) and (D.22) completes the proof. ∎
Lemma D.9 (Generalization error).
@@
$\mathrm{VCdim}({\mathscr{H}}_{t})$ $\leq cLS\log(S),$
$\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})$ $\leq cL^{2}S\log(LS),$
@@
Proof of Lemma D.9.
Since ${\mathscr{H}}_{t}\subseteq N(L,S)$, using Bartlett et al. [2019, Theorem 7] implies
(D.23) $\mathrm{VCdim}({\mathscr{H}}_{t})\leq\mathrm{VCdim}(N(L,S))\leq c_{1}LS\log(S),$
where $c_{1}$ is an absolute constant. According to Ding et al. [2025b, Lemma 13], we have $\nabla{\mathscr{H}}_{t}\subseteq N(c_{2}L,c_{3}LS)$, where $c_{2}$ and $c_{3}$ are absolute constants. Using Bartlett et al. [2019, Theorem 7] again implies
(D.24) $\mathrm{VCdim}(\nabla{\mathscr{H}}_{t})\leq c_{4}L^{2}S\log(LS),$
where $c_{4}$ is an absolute constant. This completes the proof. ∎
@@
${\mathscr{H}}_{t}\coloneq\left\{h_{t}\in N(L,S):\begin{aligned} &\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\leq\bar{B},\penalty 10000\ \inf_{\mathbf{x}\in{\mathbb{R}}^{d}}h_{t}(\mathbf{x})\geq\underline{B},\\ &\max_{1\leq k\leq d}\sup_{\mathbf{x}\in{\mathbb{R}}^{d}}|D_{k}h_{t}(\mathbf{x})|\leq 2\sigma_{T-t}^{-2}\bar{B}\end{aligned}\right\},$
where $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+8}})$. Let $\widehat{h}_{t}^{\lambda}$ be the gradient-regularized empirical risk minimizer defined as (4.6), and let $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10). Then the following inequality holds:
${\mathbb{E}}\Big[\|\nabla\log\widehat{h}_{t}^{\lambda}-\nabla\log h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}\Big]\leq C\sigma_{T-t}^{-8}n^{-\frac{2}{d+8}},$
provided that the regularization parameter $\lambda$ is set as $\lambda={\mathcal{O}}(n^{-\frac{2}{d+8}})$, where $C$ is a constant depending only on $d$, $\bar{B}$, and $\underline{B}$.
@@
Substituting Lemmas D.8, and D.9 into Lemma 5.2 yields
@@
$\leq C\frac{\log^{4}N}{\sigma_{T-t}^{8}N^{4}}+C\lambda\frac{\log^{2}N}{\sigma_{T-t}^{8}N^{2}}+C\Big(\frac{N^{d}\log^{2}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+C\frac{\lambda}{\sigma_{T-t}^{4}}\Big(\frac{N^{d}\log^{4}N}{n\log^{-1}n}\Big)^{\frac{1}{2}}+C\frac{\lambda^{2}}{\sigma_{T-t}^{8}},$
where $C$ is a constant only depending on $d$ and $\bar{B}$, and we used the fact $L\leq C^{\prime}\log N$ and $S\leq N^{d}$ in Lemma D.8. Similarly,